1. In a room full of air, the air is mainly composed of Nitrogen and Oxygen molecules (both at room temperature). Find (to twW A (97°07) = suda SW 1°xɛ Si on n oin onando CITI A ini 20 SUOISSAdy som innba
Answer #1

Vrms = ( 3kB T / m)1/2

KB is constant and equal to 1.38 x 10-23 J/K

T is tempreture in kelvin. Taking room tempreture as 27 deg C,
T = 300 K

m is mass of molecule in kg

For Oxygen , m = 5.3x10-26 kg
For nitrogen , m = 4.65x10-26 kg

For oxygen , Vrms = (3*1.38x10-23 *300/5.3x10-26)1/2 = 484 m/s
For Nitrogen, Vrms = (3*1.38x10-23 *300/4.65x10-26)1/2 = 517 m/s